# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v1

@@ 5. Experimental setup
Dense-QPP (Arabzadeh et al., 2023) is robustness-based and designed for dense retrievers only: it injects noise neural representation of the given query, and then measures the similarity between ranked lists for the original query and perturbed query representations.
@@
BERTQPP (Arabzadeh et al., 2021b) is a regression-based method, which predicts a QPP score by using BERT representations for the query and the top-ranked item. We use the cross-encoder version of BERTQPP.
@@
M-QPPF (Khodabakhsh and Bagheri, 2023) is also regression-based, which models QPP and document ranking jointly, by adopting a shared BERT layer to learn representations for query–document pairs, and using two layers to model QPP and document ranking, respectively.
QPP method | Ranker: BM25 | Ranker: ANCE | ||||||||||
TREC-DL 19 | TREC-DL 20 | TREC-DL 21 | TREC-DL 22 | TREC-DL 19 | TREC-DL 20 | |||||||
P- | K- | P- | K- | P- | K- | P- | K- | P- | K- | P- | K- | |
Clarity | 0.135 | 0.028 | 0.050 | 0.021 | 0.183 | 0.161 | 0.253∗ | 0.099 | -0.078 | -0.012 | -0.074 | -0.048 |
WIG | 0.113 | 0.164 | 0.286∗ | 0.218∗ | 0.237 | 0.206∗ | 0.029 | 0.082 | 0.313∗ | 0.228 | 0.059 | 0.048 |
NQC | 0.194 | 0.117 | 0.152 | 0.191 | 0.227 | 0.195 | 0.223 | 0.048 | 0.350∗ | 0.200 | 0.145 | 0.112 |
0.195 | 0.164 | 0.200 | 0.211∗ | 0.278∗ | 0.174 | 0.038 | 0.048 | 0.384∗ | 0.287∗ | 0.171 | 0.118 | |
n() | 0.144 | 0.181 | 0.187 | 0.123 | 0.127 | 0.140 | 0.169 | 0.113 | 0.200 | 0.176 | -0.008 | 0.022 |
SMV | 0.141 | 0.097 | 0.126 | 0.193 | 0.240 | 0.189 | 0.227∗ | 0.094 | 0.352∗ | 0.256∗ | 0.182 | 0.161 |
UEF(NQC) | 0.235 | 0.256∗ | 0.270∗ | 0.211∗ | 0.231 | 0.111 | 0.216 | 0.065 | 0.340∗ | 0.260∗ | 0.131 | 0.108 |
RLS(NQC) | 0.272 | 0.122 | 0.290∗ | 0.193 | 0.234 | 0.195 | 0.224 | 0.095 | 0.359∗ | 0.273∗ | 0.178 | 0.139 |
QPP-PRP | 0.292 | 0.189 | 0.163 | 0.184 | -0.080 | -0.017 | 0.122 | 0.091 | 0.259 | 0.246 | 0.100 | -0.008 |
Dense-QPP | - | - | - | - | - | - | - | - | 0.452∗ | 0.280∗ | 0.209 | 0.139 |
NQA-QPP | 0.181 | 0.122 | 0.062 | 0.069 | 0.161 | 0.163 | 0.224 | 0.177∗ | -0.026 | -0.009 | -0.059 | -0.080 |
BERTQPP | 0.281 | 0.136 | 0.237 | 0.155 | 0.206 | 0.134 | 0.148 | 0.122 | 0.330∗ | 0.214 | 0.046 | -0.012 |
qppBERT-PL | 0.145 | 0.138 | 0.166 | 0.152 | 0.339∗ | 0.244∗ | 0.131 | 0.206∗ | 0.092 | 0.025 | -0.224 | -0.218 |
M-QPPF | 0.317∗ | 0.208 | 0.335∗ | 0.273∗ | 0.282∗ | 0.209∗ | 0.161 | 0.187∗ | 0.292 | 0.200 | 0.068 | 0.038 |
QPP-LLM (few-shot) | 0.008 | 0.003 | -0.081 | -0.129 | -0.053 | -0.053 | -0.241 | -0.155 | -0.008 | 0.005 | -0.226 | -0.207 |
QPP-LLM (fine-tuned) | 0.171 | 0.158 | 0.228 | 0.206 | 0.030 | 0.099 | -0.038 | 0.009 | -0.073 | 0.011 | -0.022 | 0.069 |
QPP-GenRE (n=10) | 0.538†∗ | 0.486†∗ | 0.560†∗ | 0.475†∗ | 0.524†∗ | 0.435†∗ | 0.350†∗ | 0.262†∗ | 0.567†∗ | 0.440†∗ | 0.293†∗ | 0.257†∗ |
We propose two LLM-based QPP baselines. The research on using LLMs for arithmetic tasks shows that LLaMA treats numbers as distinct tokens and can understand and generate numerical values (Liu and Low, 2023). Inspired by this, we prompt LLaMA-7B to directly generate a numerical score given a query and the ranked list with passages for the query. We consider two variants:
@@
QPP evaluation and target IR measures. We follow established best practices (Datta et al., 2022c; Zamani et al., 2018; Carmel and Yom-Tov, 2010; Hauff et al., 2008; Cronen-Townsend et al., 2002) to evaluate QPP by measuring Pearson’s and Kendall’s correlation coefficients between the actual and predicted performance of a query set. We also consider scaled Mean Absolute Ranking Error (sMARE) (Faggioli et al., 2022, 2021b) and draw the same conclusion as Pearson’s and Kendall’s ; we show the results in our repository. As for target IR metrics, we consider the two most primary official IR metrics used in TREC DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019), RR@10 (precision-oriented) and nDCG@10 (recall-oriented); recent QPP studies (Arabzadeh et al., 2023; Faggioli et al., 2023e; Khodabakhsh and Bagheri, 2023) consider either or both of these metrics as their target metrics. Following (Datta et al., 2022c), we use relevance scale 2 as positive to compute actual binary IR measures (e.g., RR). When calculating correlation for nDCG@10, the actual values of nDCG@10 are calculated by human-labeled and multi-graded relevance judgments, while the nDCG@10 values predicted by QPP-GenRE are based on its generated binary relevance judgments.
@@
Implementation details. For all unsupervised QPP baselines, we tune the hyper-parameters for predicting the performance of a ranker (either BM25 or ANCE) on TREC-DL 19 (TREC-DL 21) based on Pearson’s correlation for predicting the performance of the same ranker on TREC-DL 20 (TREC-DL 22), and vice versa. We select the cutoff value for Clarity, NQC, WIG, SMV and so on from . We select hyper-parameter for n() from .
QPP method | Ranker: BM25 | Ranker: ANCE | ||||||||||
TREC-DL 19 | TREC-DL 20 | TREC-DL 21 | TREC-DL 22 | TREC-DL 19 | TREC-DL 20 | |||||||
P- | K- | P- | K- | P- | K- | P- | K- | P- | K- | P- | K- | |
Clarity | 0.091 | 0.056 | 0.358∗ | 0.250∗ | 0.137 | 0.078 | 0.202 | 0.090 | -0.088 | -0.062 | -0.091 | -0.045 |
WIG | 0.520∗ | 0.331∗ | 0.615∗ | 0.423∗ | 0.311∗ | 0.281∗ | 0.350∗ | 0.249∗ | 0.515∗ | 0.368∗ | 0.218 | 0.150 |
NQC | 0.468∗ | 0.300∗ | 0.508∗ | 0.401∗ | 0.134 | 0.221∗ | 0.360∗ | 0.156∗ | 0.548∗ | 0.372∗ | 0.411∗ | 0.290∗ |
n() | 0.532∗ | 0.311∗ | 0.622∗ | 0.443∗ | 0.328∗ | 0.234∗ | 0.336∗ | 0.228∗ | 0.388∗ | 0.315∗ | 0.103 | 0.075 |
SMV | 0.376∗ | 0.271∗ | 0.463∗ | 0.383∗ | 0.327∗ | 0.236∗ | 0.338∗ | 0.155∗ | 0.496∗ | 0.359 | 0.380∗ | 0.283∗ |
UEF(NQC) | 0.499∗ | 0.322∗ | 0.517∗ | 0.356∗ | 0.153 | 0.232∗ | 0.311∗ | 0.145 | 0.548∗ | 0.372∗ | 0.413∗ | 0.290∗ |
RLS(NQC) | 0.469∗ | 0.169 | 0.522∗ | 0.376∗ | 0.272∗ | 0.223∗ | 0.337∗ | 0.157∗ | 0.466∗ | 0.346∗ | 0.333∗ | 0.271∗ |
QPP-PRP | 0.321 | 0.181 | 0.189 | 0.157 | 0.027 | 0.004 | 0.077 | 0.012 | 0.129 | 0.049 | 0.216 | 0.121 |
Dense-QPP | - | - | - | - | - | - | - | - | 0.565∗ | 0.389∗ | 0.419∗ | 0.318∗ |
NQA-QPP | 0.210 | 0.147 | 0.244 | 0.210∗ | 0.286∗ | 0.201∗ | 0.312∗ | 0.194∗ | 0.089 | -0.038 | 0.186 | 0.113 |
BERTQPP | 0.458∗ | 0.207 | 0.426∗ | 0.300∗ | 0.351∗ | 0.223∗ | 0.369∗ | 0.229∗ | 0.222 | 0.117 | 0.137 | 0.089 |
qppBERT-PL | 0.171 | 0.175 | 0.410∗ | 0.279∗ | 0.277∗ | 0.182 | 0.300∗ | 0.242∗ | 0.116 | 0.098 | -0.119 | -0.046 |
M-QPPF | 0.404∗ | 0.254∗ | 0.435∗ | 0.297∗ | 0.265 | 0.226∗ | 0.345∗ | 0.204∗ | 0.287 | 0.160 | 0.225 | 0.177 |
QPP-LLM (few-shot) | -0.024 | -0.031 | 0.167 | 0.138 | 0.238 | 0.201 | -0.073 | -0.077 | 0.136 | 0.120 | -0.130 | -0.094 |
QPP-LLM (fine-tuned) | 0.313∗ | 0.215 | 0.309∗ | 0.254∗ | 0.264 | 0.198 | -0.075 | -0.009 | 0.203 | 0.117 | 0.081 | 0.097 |
QPP-GenRE (n=200) | 0.724†∗ | 0.474†∗ | 0.638†∗ | 0.469†∗ | 0.546†∗ | 0.435†∗ | 0.388∗ | 0.251∗ | 0.712†∗ | 0.483†∗ | 0.457†∗ | 0.343†∗ |
QPP-GenRE (n=10) | 0.605∗ | 0.482∗ | 0.490∗ | 0.323∗ | 0.462∗ | 0.350∗ | 0.316∗ | 0.245∗ | 0.624∗ | 0.406∗ | 0.306∗ | 0.238∗ |
QPP-GenRE (n=100) | 0.712∗ | 0.472∗ | 0.609∗ | 0.457∗ | 0.545∗ | 0.427∗ | 0.332∗ | 0.246∗ | 0.719∗ | 0.489∗ | 0.456∗ | 0.355∗ |
QPP-GenRE (n=1000) | 0.715∗ | 0.477∗ | 0.627∗ | 0.459∗ | 0.547∗ | 0.436∗ | 0.388∗ | 0.251 ∗ | 0.719∗ | 0.492∗ | 0.447∗ | 0.321∗ |
@@
To predict the performance of a certain ranker (either BM25 or ANCE), we train all supervised QPP baselines based on the ranked list returned by the target ranker. To predict a certain IR evaluation measure, regression-based methods (Khodabakhsh and Bagheri, 2023; Arabzadeh et al., 2021b; Hashemi et al., 2019) are trained to learn to output the target evaluation measure during training. However, our preliminary result shows that training supervised QPP baselines, especially for regression-based supervised methods (Khodabakhsh and Bagheri, 2023; Arabzadeh et al., 2021b; Hashemi et al., 2019), on the training set of MS MARCO V1 leads to inferior QPP quality for predicting the performance of ANCE. We hypothesize that this is because ANCE was originally trained on the training set of MS MARCO V1 (Xiong et al., 2021), and so the ranked list returned by ANCE on the training set of MS MARCO V1 would have higher quality than the ranked list returned by ANCE on the evaluation sets; therefore, supervised QPP methods that share the same training set with ANCE, tend to predict inflated performance on the evaluation sets, leading to degraded QPP quality. To solve the issue and ensure the consistency of the paper, we train all supervised QPP methods (including QPP-GenRE) on the development set of MS MARCO V1 for predicting the performance of either BM25 or ANCE. We train all supervised QPP methods for 5 epochs and pick the best checkpoint for predicting the performance of a ranker on TREC-DL 19 (TREC-DL 21) based on Pearson’s correlation for predicting the performance of the same ranker on TREC-DL 20 (TREC-DL 22) and vice versa. All supervised QPP baselines use bert-base-uncased,a constant learning rate (0.00002), and the Adam optimizer (Kingma and Ba, 2015).
@@
For QPP-LLM, we prompt LLaMA-7B with the top-k retrieved items, where sets to 10. For QPP-LLM (few-shot), we randomly sample demonstration examples from the development set of MS MARCO V1; our preliminary experiments show that sampling 4 demonstrations works best. For QPP-LLM (fine-tuned), we fine-tune LLaMA-7B using PEFT as QPP-GenRE fine-tunes LLMs.
@@
We equip QPP-GenRE with LLaMA-7B for judging relevance. We use the novel PEFT method, 4-bit QLoRA (Dettmers et al., 2023), to fine-tune LLaMA-7B. The training of judging relevance needs positive and negative items per query. For positive items, we use the items annotated as relevant in qrels; we randomly sample one negative item from the ranked list (1000 items) returned by BM25 per query. we fine-tune LLaMA-7B for 5 epochs on the development set of MS MARCO V1, taking about an hour and a half on an NVIDIA A100 GPU (40GB).
@@ 6. Results
Predicting a precision-oriented IR measure. To answer RQ1, we compare QPP-GenRE and all baselines in predicting the performance of BM25 and ANCE w.r.t. a widely-used precision-oriented metric, RR@10; see Table 1. We have three main observations.
First, our proposed method, QPP-GenRE, outperforms all the baselines in terms of all correlation coefficients on all datasets when predicting the performance of both rankers. In particular, we observe that QPP-GenRE outperforms QPP-PRP (Singh et al., 2023) which is a recently proposed baseline by 84% (0.292 vs. 0.538) in terms of Pearson’s when predicting RR@10 for BM25 on TREC-DL 19.
@@
Third, there is no clear winner among the baselines, and the performance of baselines shows a bigger variance than QPP-GenRE across different datasets and rankers. E.g., the unsupervised method WIG achieves a good result among baselines for assessing BM25 on TREC-DL 20, while it gets nearly zero correlation coefficients on TREC-DL 22 when assessing BM25. Conversely, QPP-GenRE consistently achieves the best performance across different datasets, rankers, showcasing robust performance.
Predicting a recall-oriented IR measure. To answer RQ2, Table 2 lists the performance of QPP-GenRE along with all the baselines on assessing BM25 and ANCE in terms of nDCG@10, a widely-used recall-oriented metric. For QPP-GenRE, we universally set the judging depth as 200 for all evaluation sets. The result reveals that by judging only 200 items per query, we can achieve state-of-the-art QPP quality in terms of nDCG@10 for all rankers on all evaluation sets; we will go into the impact of judging depth on QPP-GenRE’s performance in the next section. Also, QPP-LLM (few-shot) and QPP-LLM (fine-tuning) are among the worst-performing baselines, showing that the LLM struggles to generate numerical scores.

### v2

@@ 5.4. QPP baselines
Dense-QPP (Arabzadeh et al., 2023) is robustness-based and designed for dense retrievers only: it injects noise neural representation of the given query, and then measures the similarity between ranked lists for the original query and perturbed query representations. Note that Dense-QPP (Arabzadeh et al., 2023) is designed for predicting the ranking quality of neural-based retrievers; it cannot predict the ranking quality of BM25.
@@
BERTQPP (Arabzadeh et al., 2021b) is a regression-based method, which predicts a QPP score by using BERT representations for the query and the top-ranked item. We use the cross-encoder version of BERTQPP because of its promising results.
@@
M-QPPF (Khodabakhsh and Bagheri, 2023) is also regression-based and models QPP and document ranking jointly, by adopting a shared BERT layer to learn representations for query-document pairs, and using two layers to model QPP and document ranking, respectively.
While to the best of our knowledge there is no LLM-based QPP method yet, to have a fair comparison with LLM-based approaches, we propose two LLM-based QPP baselines. Research on using LLMs for arithmetic tasks shows that LLaMA treats numbers as distinct tokens and can understand and generate numerical values (Liu and Low, 2023). Inspired by this, we prompt LLaMA-7B to directly generate a numerical score given a query and the ranked list with passages for the query; the prompt is shown in Figure 2. We consider two variants:
@@ 5.5. QPP evaluation and target IR evaluation measures
We follow established best practices (Datta et al., 2022c; Zamani et al., 2018; Carmel and Yom-Tov, 2010; Hauff et al., 2008; Cronen-Townsend et al., 2002) to evaluate QPP by measuring linear correlation by Pearson’s as well as ranked-based correlation through Kendall’s correlation coefficients between the actual and predicted performance of a query set.444We also consider scaled Mean Absolute Ranking Error (sMARE) (Faggioli et al., 2022, 2021b) and draw the same conclusion as Pearson’s and Kendall’s ; we include the results in our repository at https://github.com/ChuanMeng/QPP-GenRE. As for target IR metrics, we consider the two primary official IR metrics used in TREC DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019), RR@10 (precision-oriented) and nDCG@10 (considering recall); recent QPP studies (Arabzadeh et al., 2023; Faggioli et al., 2023e; Khodabakhsh and Bagheri, 2023) consider either or both of these metrics as their target metrics. Following (Datta et al., 2022c), we use relevance scale 2 as positive to compute actual binary IR measures (e.g., RR). When calculating correlation for nDCG@10, the actual values of nDCG@10 are calculated by human-labeled and multi-graded relevance judgments, while the nDCG@10 values predicted by QPP-GenRE are based on its generated binary judgments.
5.6. Implementation details
@@
For all unsupervised QPP baselines, we tune the hyper-parameters for predicting the ranking quality of a ranker (either BM25 or ANCE) on TREC-DL 19 (TREC-DL 21) based on Pearson’s correlation for predicting the ranking quality of the same ranker on TREC-DL 20 (TREC-DL 22), and vice versa. We select the cut-off value for Clarity, NQC, WIG, SMV and so on from . n() has a a hyper-parameter , which we choose from the set .
@@
To predict the performance of a certain ranker (either BM25 or ANCE), we train all supervised QPP baselines based on the ranked list returned by the target ranker. To predict a certain IR evaluation measure, regression-based methods (Khodabakhsh and Bagheri, 2023; Arabzadeh et al., 2021b; Hashemi et al., 2019) are trained to learn to output the target evaluation measure during training. However, our preliminary result shows that training supervised QPP baselines, especially for regression-based supervised methods (Khodabakhsh and Bagheri, 2023; Arabzadeh et al., 2021b; Hashemi et al., 2019), on the training set of MS MARCO V1 leads to inferior QPP quality for predicting the performance of ANCE. We hypothesize that this is because ANCE was originally trained on the training set of MS MARCO V1 (Xiong et al., 2021), and so the ranked list returned by ANCE on the training set of MS MARCO V1 would have higher quality than the ranked list returned by ANCE on the evaluation sets; therefore, supervised QPP methods that share the same training set as ANCE, tend to predict inflated performance on the evaluation sets, leading to degraded QPP quality. To solve the issue and ensure the consistency of the paper, we train all supervised QPP methods (including QPP-GenRE) on the development set of MS MARCO V1 (6980 queries) for predicting the performance of either BM25 or ANCE. We train all supervised QPP methods for 5 epochs and pick the best checkpoint for predicting the performance of a ranker on TREC-DL 19 (TREC-DL 21) based on Pearson’s correlation for predicting the performance of the same ranker on TREC-DL 20 (TREC-DL 22) and vice versa. All supervised QPP baselines use bert-base-uncased,555https://github.com/huggingface/transformers a constant learning rate (0.00002), and the Adam optimizer (Kingma and Ba, 2015).
@@
For QPP-LLM, we prompt LLaMA-7B with the top- retrieved items, where is set to 10. For QPP-LLM (few-shot), we randomly sample demonstration examples from the development set of MS MARCO V1; our preliminary experiments show that sampling 2 demonstrations works best. For QPP-LLM (fine-tuned), we fine-tune LLaMA-7B using PEFT as QPP-GenRE fine-tunes LLMs.
@@
We equip QPP-GenRE with LLaMA-7B for judging relevance. We use a recent PEFT method, 4-bit QLoRA (Dettmers et al., 2023), to fine-tune LLaMA-7B. The training of judging relevance needs positive and negative items per query. For positive items, we use the items annotated as relevant in qrels per query; we randomly sample one negative item from the ranked list (1,000 items) returned by BM25 per query. We fine-tune LLaMA-7B for 5 epochs on the development set of MS MARCO V1, taking about an hour and a half on an NVIDIA A100 GPU (40GB).
QPP method | Ranker: BM25 | |||||||
TREC-DL 19 | TREC-DL 20 | TREC-DL 21 | TREC-DL 22 | |||||
P- | K- | P- | K- | P- | K- | P- | K- | |
Clarity | 0.135 | 0.028 | 0.050 | 0.021 | 0.183 | 0.161 | 0.253∗ | 0.099 |
WIG | 0.113 | 0.164 | 0.286∗ | 0.218∗ | 0.237 | 0.206∗ | 0.029 | 0.082 |
NQC | 0.194 | 0.117 | 0.152 | 0.191 | 0.227 | 0.195 | 0.223 | 0.048 |
0.195 | 0.164 | 0.200 | 0.211∗ | 0.278∗ | 0.174 | 0.038 | 0.048 | |
n() | 0.144 | 0.181 | 0.187 | 0.123 | 0.127 | 0.140 | 0.169 | 0.113 |
SMV | 0.141 | 0.097 | 0.126 | 0.193 | 0.240 | 0.189 | 0.227∗ | 0.094 |
UEF(NQC) | 0.235 | 0.256∗ | 0.270∗ | 0.211∗ | 0.231 | 0.111 | 0.216 | 0.065 |
RLS(NQC) | 0.272 | 0.122 | 0.290∗ | 0.193 | 0.234 | 0.195 | 0.224 | 0.095 |
QPP-PRP | 0.292 | 0.189 | 0.163 | 0.184 | -0.080 | -0.017 | 0.122 | 0.091 |
NQA-QPP | 0.181 | 0.122 | 0.062 | 0.069 | 0.161 | 0.163 | 0.224 | 0.177∗ |
BERTQPP | 0.281 | 0.136 | 0.237 | 0.155 | 0.206 | 0.134 | 0.148 | 0.122 |
qppBERT-PL | 0.145 | 0.138 | 0.166 | 0.152 | 0.339∗ | 0.244∗ | 0.131 | 0.206∗ |
M-QPPF | 0.317∗ | 0.208 | 0.335∗ | 0.273∗ | 0.282∗ | 0.209∗ | 0.161 | 0.187∗ |
QPP-LLM (few-shot) | 0.008 | 0.003 | -0.081 | -0.129 | -0.053 | -0.053 | -0.241 | -0.155 |
QPP-LLM (fine-tuned) | 0.171 | 0.158 | 0.228 | 0.206 | 0.030 | 0.099 | -0.038 | 0.009 |
QPP-GenRE () | 0.538†∗ | 0.486†∗ | 0.560†∗ | 0.475†∗ | 0.524†∗ | 0.435†∗ | 0.350†∗ | 0.262†∗ |
QPP method | Ranker: ANCE | |||
TREC-DL 19 | TREC-DL 20 | |||
P- | K- | P- | K- | |
Clarity | -0.078 | -0.012 | -0.074 | -0.048 |
WIG | 0.313∗ | 0.228 | 0.059 | 0.048 |
NQC | 0.350∗ | 0.200 | 0.145 | 0.112 |
0.384∗ | 0.287∗ | 0.171 | 0.118 | |
n() | 0.200 | 0.176 | -0.008 | 0.022 |
SMV | 0.352∗ | 0.256∗ | 0.182 | 0.161 |
UEF(NQC) | 0.340∗ | 0.260∗ | 0.131 | 0.108 |
RLS(NQC) | 0.359∗ | 0.273∗ | 0.178 | 0.139 |
QPP-PRP | 0.259 | 0.246 | 0.100 | -0.008 |
Dense-QPP | 0.452∗ | 0.280∗ | 0.209 | 0.139 |
NQA-QPP | -0.026 | -0.009 | -0.059 | -0.080 |
BERTQPP | 0.330∗ | 0.214 | 0.046 | -0.012 |
qppBERT-PL | 0.092 | 0.025 | -0.224 | -0.218 |
M-QPPF | 0.292 | 0.200 | 0.068 | 0.038 |
QPP-LLM (few-shot) | -0.008 | 0.005 | -0.226 | -0.207 |
QPP-LLM (fine-tuned) | -0.073 | 0.011 | -0.022 | 0.069 |
QPP-GenRE () | 0.567†∗ | 0.440†∗ | 0.293†∗ | 0.257†∗ |
@@ 6. Results
6.1. Predicting a precision-oriented IR measure
To answer RQ1, we compare QPP-GenRE and all baselines in predicting the performance of BM25 and ANCE w.r.t. a widely-used precision-oriented metric, RR@10; see Table 2. We have three main observations.
First, our proposed method, QPP-GenRE, outperforms all baselines in terms of both correlation coefficients on all datasets when predicting the performance of both rankers. In particular, we observe that QPP-GenRE outperforms QPP-PRP (Singh et al., 2023), which is a recently proposed baseline by 84% (0.292 vs. 0.538) in terms of Pearson’s when predicting RR@10 for BM25 on TREC-DL 19.
@@ 6.1. Predicting a precision-oriented IR measure
Third, there is no clear winner among the baselines, and the performance of baselines shows a bigger variance than QPP-GenRE across different datasets and rankers. E.g., the unsupervised method WIG achieves a good result among baselines for assessing BM25 on TREC-DL 20, while it gets nearly zero correlation coefficients on TREC-DL 22 when assessing BM25. Conversely, QPP-GenRE consistently achieves the best performance across datasets and rankers, thus showing robust performance.
QPP method | TREC-DL 19 | TREC-DL 20 | TREC-DL 21 | TREC-DL 22 | ||||
P- | K- | P- | K- | P- | K- | P- | K- | |
Clarity | 0.091 | 0.056 | 0.358∗ | 0.250∗ | 0.137 | 0.078 | 0.202 | 0.090 |
WIG | 0.520∗ | 0.331∗ | 0.615∗ | 0.423∗ | 0.311∗ | 0.281∗ | 0.350∗ | 0.249∗ |
NQC | 0.468∗ | 0.300∗ | 0.508∗ | 0.401∗ | 0.134 | 0.221∗ | 0.360∗ | 0.156∗ |
0.478∗ | 0.327∗ | 0.529∗ | 0.440∗ | 0.298∗ | 0.258∗ | 0.142∗ | 0.196∗ | |
n() | 0.532∗ | 0.311∗ | 0.622∗ | 0.443∗ | 0.328∗ | 0.234∗ | 0.336∗ | 0.228∗ |
SMV | 0.376∗ | 0.271∗ | 0.463∗ | 0.383∗ | 0.327∗ | 0.236∗ | 0.338∗ | 0.155∗ |
UEF(NQC) | 0.499∗ | 0.322∗ | 0.517∗ | 0.356∗ | 0.153 | 0.232∗ | 0.311∗ | 0.145 |
RLS(NQC) | 0.469∗ | 0.169 | 0.522∗ | 0.376∗ | 0.272∗ | 0.223∗ | 0.337∗ | 0.157∗ |
QPP-PRP | 0.321 | 0.181 | 0.189 | 0.157 | 0.027 | 0.004 | 0.077 | 0.012 |
NQA-QPP | 0.210 | 0.147 | 0.244 | 0.210∗ | 0.286∗ | 0.201∗ | 0.312∗ | 0.194∗ |
BERTQPP | 0.458∗ | 0.207 | 0.426∗ | 0.300∗ | 0.351∗ | 0.223∗ | 0.369∗ | 0.229∗ |
qppBERT-PL | 0.171 | 0.175 | 0.410∗ | 0.279∗ | 0.277∗ | 0.182 | 0.300∗ | 0.242∗ |
M-QPPF | 0.404∗ | 0.254∗ | 0.435∗ | 0.297∗ | 0.265 | 0.226∗ | 0.345∗ | 0.204∗ |
QPP-LLM (few-shot) | -0.024 | -0.031 | 0.167 | 0.138 | 0.238 | 0.201 | -0.073 | -0.077 |
QPP-LLM (fine-tuned) | 0.313∗ | 0.215 | 0.309∗ | 0.254∗ | 0.264 | 0.198 | -0.075 | -0.009 |
QPP-GenRE () | 0.724†∗ | 0.474†∗ | 0.638†∗ | 0.469†∗ | 0.546†∗ | 0.435†∗ | 0.388∗ | 0.251∗ |
QPP-GenRE () | 0.605∗ | 0.482∗ | 0.490∗ | 0.323∗ | 0.462∗ | 0.350∗ | 0.316∗ | 0.245∗ |
QPP-GenRE () | 0.712∗ | 0.472∗ | 0.609∗ | 0.457∗ | 0.545∗ | 0.427∗ | 0.332∗ | 0.246∗ |
QPP-GenRE () | 0.715∗ | 0.477∗ | 0.627∗ | 0.459∗ | 0.547∗ | 0.436∗ | 0.388∗ | 0.251∗ |
QPP method | TREC-DL 19 | TREC-DL 20 | ||
P- | K- | P- | K- | |
Clarity | -0.088 | -0.062 | -0.091 | -0.045 |
WIG | 0.515∗ | 0.368∗ | 0.218 | 0.150 |
NQC | 0.548∗ | 0.372∗ | 0.411∗ | 0.290∗ |
0.455∗ | 0.339∗ | 0.403∗ | 0.288∗ | |
n() | 0.388∗ | 0.315∗ | 0.103 | 0.075 |
SMV | 0.496∗ | 0.359 | 0.380∗ | 0.283∗ |
UEF(NQC) | 0.548∗ | 0.372∗ | 0.413∗ | 0.290∗ |
RLS(NQC) | 0.466∗ | 0.346∗ | 0.333∗ | 0.271∗ |
QPP-PRP | 0.129 | 0.049 | 0.216 | 0.121 |
Dense-QPP | 0.565∗ | 0.389∗ | 0.419∗ | 0.318∗ |
NQA-QPP | 0.089 | -0.038 | 0.186 | 0.113 |
BERTQPP | 0.222 | 0.117 | 0.137 | 0.089 |
qppBERT-PL | 0.116 | 0.098 | -0.119 | -0.046 |
M-QPPF | 0.287 | 0.160 | 0.225 | 0.177 |
QPP-LLM (few-shot) | 0.136 | 0.120 | -0.130 | -0.094 |
QPP-LLM (fine-tuned) | 0.203 | 0.117 | 0.081 | 0.097 |
QPP-GenRE () | 0.712†∗ | 0.483†∗ | 0.457†∗ | 0.343†∗ |
QPP-GenRE () | 0.624∗ | 0.406∗ | 0.306∗ | 0.238∗ |
QPP-GenRE () | 0.719∗ | 0.489∗ | 0.456∗ | 0.355∗ |
QPP-GenRE () | 0.719∗ | 0.492∗ | 0.447∗ | 0.321∗ |
6.2. Predicting an IR measure considering recall
To answer RQ2, Table 4 lists the performance of QPP-GenRE along with all the baselines on assessing BM25 and ANCE in terms of nDCG@10. For QPP-GenRE, we universally set the judging depth to 200 for all evaluation sets. The result reveals that by judging only 200 items per query, we can achieve state-of-the-art QPP quality in terms of nDCG@10 for all rankers on all evaluation sets; we will investigate the impact of judging depth on QPP-GenRE’s performance in the next section. Also, QPP-LLM (few-shot) and QPP-LLM (fine-tuning) are among the worst-performing baselines, showing that the LLMs struggle to generate numerical scores. Different from the results for RQ1, most QPP methods tend to perform better when predicting nDCG@10 than RR@10; this observation indicates that predicting RR@10 is a more challenging task.